home studios equipment staff/friends booking/rates for sale forum contact

Electrical's Old Repair List

All your geeky questions about electronics, microphones, tape machines, etc.

Moderators: kerble, Electrical-Staff

Postby greg on Mon Apr 21, 2008 2:19 pm

Equipment: UREI LA-12 ch1
In: bypass relay "facked" somehow influenced circuit
Out:
Repaired: replaced relay with an American Zetler relay (modified coil current for it)
Date: 4/01

Equipment: Ampex ATR102 audio card (-16 verision, grease fire)
In: noise visible on VU meter in repro
Out:
Repaired: Q1 noisy, replaced
Date: 4/16

Equipment: Ampex ATR102 audio card
In: distortion found when tapping main audio PWA during repro 'play' and poor LF playback response
Out:
Repaired: leaky, sputtery tantalum cap C6
Date: 4/20

Equipment: Lomo 19.9 blue
In: low op level, dirty capsule
Out:
Repaired:
Date: 4/24
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 2:22 pm

Equipment: Neotek Series 2 CH 25
In: solo/mute logic not working
Out:
Repaired: bad DP3T switch replaced (part of our mute/solo grouping mod)
Date: 3/03

Equipment: Neumann 582 PSU
In: low heater V
Out:
Repaired: rebuilt that part of the PSU
Date: 5/01

Equipment: ATR #3
In: Paper leader riding up take up tension roller.
Out:
Repaired: Improper mounting of take up tension roller
Date: 5/01
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 2:24 pm

Equipment: Garnet BTO
In: pop, pop pop noise
Out:
Repaired:replaced the neon lamp on the 'bass boost' switch
Date: 5/01

Equipment: Lomo 19-13 mic cable
In: No output
Out:
Repaired: broken wire in cable repaired
Date: 6/01

Equipment: Altec 150 s/n 4254
In: sputtering noise
Out:
Repaired: Tube replaced
Date: 6/01

Equipment: Altec 150 s/n813
In: new ebay mic
Out:
Repaired: ferrite beaded and ready to test
Date: 6/01
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 2:38 pm

Equipment: TC 2290
In: Intermittent input volume.
Out: to TC Electronics for total refurb.
Repaired:
Date: 7/01
Turns out, the input is routed through several normally closed contacts on isolated tin 1/4'' TRS connectors for loop send/return, parallel ins. Look at this shit.
Image The connections, most of them never used, oxidize and undermine the unit. Thankfully they aren't pricks about it, and will upgrade the box with gold plated connectors on the cheap.

Equipment: AKC CK1 capsule SN25063
In: ripped membrane
Out: to AKG
Repaired: $80 repair
Date: 7/01

Equipment: Sontec EQ
In: noise on return ch.
Out:
Repaired: recapped entire unit (electrolytic + tantalums), and replaced noisy 5534s problem ended up being horrible discrete op amps
Date: 8/01

Equipment: UREI LA-12
In: 'auto' always on
Out:
Repaired: Found cold/intermitent solder joints on ratio/auto pot-switch. Re-soldered all pots and switches, moved link/dual switch from the rear to front face plate
Date: 8/01
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:36 pm

Equipment: GML Preamp
In:
Out:
Repaired: Moved global phantom power switch to the front pannel
Date: 8/01

Equipment: Manley CR-3 SN
In: noise
Out:
Repaired: replaced fet
Date: 8/01

Equipment: Ampex 351 mic pre #1
In: fuses blow
Out:
Repaired: bad rectifier tube
Date: 8/01
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:38 pm

Equipment: Sovtek MIG50
In: loose parts
Out:
Repaired: replaced screws and knobs
Date: 10/01

Date: 10/01
Equipment: Earthworks Z30X
SN: 5082
Problem: oscillation on output
Out to: earthworks
Repaired: no explanation
Date: 11/1

Date: 10/01
Equipment: josephson 609 w/ KA32
SN: 2744 / 2722
Problem: noise
Out to: josephson
Repaired: broken solder connection
Date: 11/13/01

Date: 11/5/01
Equipment: ATR102 4 speed padnets (grease fire versions)
SN:
Problem: callibration problems in trim pots
Out to:
Repaired: replaced all trim pots related to 30/15ips operation
Date: 10/5/01
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:43 pm

Date: 11/9/01
Equipment: UREI/electrical EQ
SN: 2
Problem: checked out
Out to:
Repaired: construction complete
Date: 1/5

Date: 11/13/01
Equipment: GML 100 crown
SN:
Problem: no output
Out to:
Repaired: broken wires + lost set screw in XLR housing (lots of stress on wire)
Date: 11/13/01

Date: 11/13/01
Equipment: ATR #4 set up for 4 speeds
SN:
Problem:
Out to:
Repaired:
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:44 pm

Date: 11/13/01
Equipment: MX.35 #2
SN: 1030
Problem: No o/p on mic chan 2
Out to:
Repaired: R8 47K 1W resistor opened up
Date:

Date: 11/15/01
Equipment: Neotek Ser. II
SN:
Problem: no o/p on echo return 3+4
Out to:
Repaired: faulty DPDT switch '2-mix' send - replaced
Date: 11/16/01

Date: 12/07/01
Equipment: Neotek Ser. II
SN:
Problem: CH1 Fader Sticky
Out to:
Repaired: lubed it up and switched it w/ ch. 36
Date: 12/07/01

Date: 11/01
Equipment: STC (BBC) 4038
SN:
Problem: Flattened ribbon in middle
Out to:
Repaired: redone by Coles
Date: 8/02
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:48 pm

Date: 12/01
Equipment: ATR #2 (ATR serviced)
SN:
Problem: tape-lifter sleeves cut sharp - sheds tape - can't be turned
Out to:
Repaired: With the headblock out, I carefully heated up the lifters with a heat gun and pulled the sleeves out, then reset them with unworn side out.
Date:

Date: 12/07/01
Equipment: Tension adjust
SN: 820 STB
Problem:
Out to:
Repaired:
Date:

Date: 12/09/01
Equipment: 451 body
SN: 25515
Problem: no o/p
Out to: AKG $60 repair
Repaired: see service sheets in file cabinet
Date: 1/21/02

Date: 1/02/02
Equipment: Sontec EQ
SN:
Problem: noise in L ch.
Out to:
Repaired:
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:52 pm

Date: 1/08/02
Equipment: MCI JH110 B 2TK
SN: 3176
Problem: unreliable as a whole
Out to:
Repaired: Went through machine completely
Date:

Date: 1/08/02
Equipment: Neumann 582
SN: 3115
Problem: wind/blowing noise
Out to:
Repaired: cleaned and checked tube socket, capsule conn.
Date:

Date: 1/09/02
Equipment: Neotek II ch. 25
SN:
Problem: no o/p on direct out or buss send with 'mixdown' in
Out to:
Repaired: broken jumper fixed
Date: 1/08/2

Date: 1/09/02
Equipment: Neotek II mute group lamp
SN:
Problem: broken PSU wire
Out to:
Repaired: fixed
Date: 1/10/02
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:53 pm

Date: 1/10/02
Equipment: 2X SM98 Mic Cables
SN:
Problem: damaged
Out to:
Repaired: fixed
Date: 1/10/02

Date: 1/9/02
Equipment: Studer A820
SN: blackside
Problem:
Out to:
Repaired: lubed capstan
Date: 1/09/02

Date: 1/11/02
Equipment: Studer A820
SN: brownside
Problem:
Out to:
Repaired: lubed capstan
Date: 1/11/02

Date: 1/14/02
Equipment: SR'A'SSER CM 060
SN: 6413
Problem: Americanizing power + audio conn. on PSU
Out to:
Repaired: mic works in quick test
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 3:58 pm

Date: 1/15/02
Equipment: Studer (black)
SN: 10XX
Problem:
Out to:
Repaired: calibrated PS Rails and Corrected tack pulse PWM
Date:

Date: 1/15/02
Equipment: Neotek Series II
SN:
Problem: No mic in on channel 16
Out to:
Repaired: Short from GND to positive discovered only when the mic line to console edac was connected. The shied from another channel was shorting against channel 16's pin in the edac connector
Date: 1/15/02

Date: 1/18/02
Equipment: ATR Transport control card
SN: spare
Problem:
Out to:
Repaired: replaced 2 broken .1u caps
Date:

Date: 1/21/02
Equipment: Furman HP cable
SN:
Problem: broken crimp
Out to:
Repaired: replaced conn.
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:00 pm

Date: 1/22/02
Equipment: 351 studio B 1:2
SN:
Problem: polarity reversed
Out to:
Repaired: fixed on polarity switch
Date:

Date: 1/24/02
Equipment: headphones
SN:
Problem: no left speaker
Out to:
Repaired: replaced 1/4'' conn
Date:

Date: 1/24/02
Equipment: Manley CR-3A(B)
SN: #1 on transformer
Problem: noise
Out to:
Repaired: zener diode (12V) failed
Date: 1/24/02

Date: 1/28/02
Equipment: Orang Lo Stressor
SN:
Problem: EQ not working
Out to: Shea Ako repaired
Repaired:
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:05 pm

Date: 1/30/02
Equipment: Neotek II CH24
SN: 27
Problem: send to Left/Right mix intermittent do to header connector
Out to:
Repaired: replaced w/ gold plated header for testing
Date: 1/30/02

Date: 1/30/02
Equipment: Neotek II CH23
SN: 26
Problem: PFL Not Working
Out to:
Repaired: no info
Date:

Date: 1/30/02
Equipment: Studer 820 Blacksides high frequency divider boards chan 17+19
SN:
Problem: CH17+19 squeel on spot erase, squeal recorded from the erase head
Out to:
Repaired: replaced C25 and C18 problem went away. C25 possibly a bypass cap for IC 15. If spurious oscillations occurs there it might make it to the erase head ??
Date: 1/31/02

Date: 2/1/02
Equipment: Eventide H3000
SN: 5148?
Problem: noise on samples
Out to: Eventide again 2/6, last time 9/01
Repaired:
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:11 pm

Date: 2/6/01
Equipment: Garnet BTO
SN: 5350
Problem: LFO out of transients, pops every 10 seconds
Out to:
Repaired: changed cap going from 65N7's Plate to o/p diver section to .5uF 600V - neon lamp replaced w/ LED (Radio Shack 276-351)
Date: 2/7/02

Date: 2/6/02
Equipment: ATR headstack
SN: 1/2 #2
Problem: squeeling scrape fultter idler
Out to: JRF
Repaired: relap - relube
Date: 2/13/02

Date: 2/9/02
Equipment: Traynor PA cabinets 2X1" tweeter drivers, 1 15" speaker
SN:
Problem: busted
Out to: J&J repair
Repaired: check
Date: 2/18/02
Last edited by greg on Mon Apr 21, 2008 4:19 pm, edited 1 time in total.
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:13 pm

Date: 2/11/02
Equipment: Garnet BTO
SN: 1144690
Problem: popping and ticking noise
Out to:
Repaired: neon lamps were somehow infering w/ B+ of 12AX7's replaced w/ LED's
Date: 2/11/02

Date: 2/12/02
Equipment: Soundfield MKIV
SN: 517
Problem: sputtering noise
Out to: soundfield
Repaired: fixed and calibrated
Date:

Date: 2/12/02
Equipment: Traynor T55OB
SN:
Problem:
Out to:
Repaired: ok, nothing
Date: 2/17/02

Date: 2//12/02
Equipment: Sovtek MIG50H
SN:
Problem: Blown power transformer and resistors and tubes
Out to:
Repaired: all parts replaced
Date: 2/28/02
Old Transformer looked charred. I think I still have it.
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:19 pm

Date: 2/19/02
Equipment: Lomo 19-18
SN: 690018
Problem: noise and hum and popping,
Out to:
Repaired: changed tube (5703) changed fil voltage on PSU to 5.9v (from 4v)

Date: 2/19/02
Equipment: Lomo 19-18 capsule
SN: 680048
Problem: low o/p -6 dB compared to other 19-18
Out to:
Repaired: cleaned capsule connectors, not as dramatic as before (-2dB)

Date: 1/28/02
Equipment: AKG C-12 PSU
SN:
Problem: broken power switch
Out to:
Repaired: replaced
Date: 2/21/02

Date: 1/21/02
Equipment: Sennheiser 421
SN: bakelite albino vers.
Problem: buzz on o/p
Out to:
Repaired: Rewired connector
Date:

Date: 1/22/02
Equipment: Sontec 205B
SN: 209
Problem: noise and refurbishment
Out to:
Repaired: calibrated noiseles sw replaced discrete op amp + 2 others
Date: 2/26/02
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:21 pm

Date: 2/23/02
Equipment: Orange Amp OR120
SN: 3525
Problem: cutting out + noise
Out to:
Repaired:
Date:

Date: 3/08/02
Equipment: Josephson
SN: KA32 Cap2726
Problem: rattle in capsule
Out to:
Repaired: removed chunk of solder from capsule housing
Date: 3/4/02

Date: 3/18/02
Equipment: Studer (new one)
SN: 1415
Problem: eprom software update
Out to:
Repaired: now up to 4/91 software
Date:

Date: 3/18/02
Equipment: milleniamedia
SN:
Problem:
Out to:
Repaired: tube mod telefunken
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:32 pm

Date: 3/XX/02
Equipment: Neotek Elite?
SN:
Problem: no line in on channel 2
Out to:
Repaired: re-punched wires on patch bay
Date:

Date: 3/xx/02
Equipment: Studio A 24 track "b" patch bay wiring
SN:
Problem: too many wires punched on the same points causing loose, intermittent connections.
Out to:
Repaired: moved parallel 24-track "b" output wiring (from tie line panel in the control room) to parallel connection points on the "line in" row of the patch bay. Now, there are no more than 2 wires punched on any given post.
Date:

Date: 3/26/02
Equipment: ATR #3 VSO
SN: XX1XX..
Problem:
Out to: taken to ATR service for complete overhaul
Repaired:
Date:

Date: 4/3/02
Equipment: Neotek Elite
SN: CH13
Problem: signal going to aux B when level is down
Out to:
Repaired: replaced bad pot (dual 100k pot),
Also reversed a meter mod, which had the meter see the line in, pre line trim, because that is how the rest of the console is.
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

Postby greg on Mon Apr 21, 2008 4:40 pm

Date: 4/4/02
Equipment: Lomo 19-9
SN: 331 009
Problem: dirty capsules - corrosion
Out to:
Repaired:
Date:

Date:
Equipment: MCI 110 2TK
SN:
Problem: left chan. record meter back and forth
Out to:
Repaired:
Date:

Date: 4/4/02
Equipment: Studer (blacksides) 24TK Internal NRS Cards
SN:
Problem: Noise reduction not sounding right on some channels
Out to:
Repaired: out of spec bipolar caps rolling low frequencies off in different modes- replaced w/ xicon BP caps. all boards to be recapped.
Date: 4/8/02

Date: 5/24/02
Equipment: Neotek 'elite'
SN:
Problem: speaker select switch C scratchy and intermittent.
Out to:
Repaired: replaced speaker select switches A,B, and C with gold plated equivalent.
Date:
User avatar
greg
Master Of The Computer
Master Of The Computer
 
Posts: 4935
Joined: Wed May 14, 2003 7:26 pm
Location: Chicago

PreviousNext

Return to Tech Room

Who is online

Users browsing this forum: No registered users and 4 guests